# Language Models as Black-Box Optimizers for Vision-Language Models

### v4

@@ 10 T2I Experimental Details
![[Uncaptioned image]](extracted/5581488/t2i/orange_original.png)
![[Uncaptioned image]](extracted/5581488/t2i/orange_initial.png)
@@
![[Uncaptioned image]](extracted/5581488/t2i/knight_original.png)
![[Uncaptioned image]](extracted/5581488/t2i/knight_initial.png)
@@
![[Uncaptioned image]](extracted/5581488/t2i/dove_original.png)
![[Uncaptioned image]](extracted/5581488/t2i/dove_initial.png)
@@
![[Uncaptioned image]](extracted/5581488/t2i/dino_original.png)
![[Uncaptioned image]](extracted/5581488/t2i/dino_initial.png)
@@
![[Uncaptioned image]](extracted/5581488/t2i/show_example3_original.jpg)
![[Uncaptioned image]](extracted/5581488/t2i/show_example3_init.jpg)
@@
![[Uncaptioned image]](extracted/5581488/t2i/show_example5_original.jpg)
![[Uncaptioned image]](extracted/5581488/t2i/show_example5_init.jpg)
@@
![[Uncaptioned image]](extracted/5581488/t2i/failure_art_original.png)
@@
![[Uncaptioned image]](extracted/5581488/t2i/failure_count_original.png)
@@
![[Uncaptioned image]](extracted/5581488/t2i/failure_cube_original.png)
@@
![[Uncaptioned image]](extracted/5581488/t2i/failure_panels_original.png)
@@
![[Uncaptioned image]](extracted/5581488/t2i/japan_original.png)
![[Uncaptioned image]](extracted/5581488/t2i/japan_initial.png)
@@
![[Uncaptioned image]](extracted/5581488/t2i/panda_original.png)
![[Uncaptioned image]](extracted/5581488/t2i/panda_initial.png)

### v5

@@ 10 T2I Experimental Details
![[Uncaptioned image]](extracted/2309.05950v5/t2i/orange_original.png)
![[Uncaptioned image]](extracted/2309.05950v5/t2i/orange_initial.png)
@@
![[Uncaptioned image]](extracted/2309.05950v5/t2i/knight_original.png)
![[Uncaptioned image]](extracted/2309.05950v5/t2i/knight_initial.png)
@@
![[Uncaptioned image]](extracted/2309.05950v5/t2i/dove_original.png)
![[Uncaptioned image]](extracted/2309.05950v5/t2i/dove_initial.png)
@@
![[Uncaptioned image]](extracted/2309.05950v5/t2i/dino_original.png)
![[Uncaptioned image]](extracted/2309.05950v5/t2i/dino_initial.png)
@@
![[Uncaptioned image]](extracted/2309.05950v5/t2i/show_example3_original.jpg)
![[Uncaptioned image]](extracted/2309.05950v5/t2i/show_example3_init.jpg)
@@
![[Uncaptioned image]](extracted/2309.05950v5/t2i/show_example5_original.jpg)
![[Uncaptioned image]](extracted/2309.05950v5/t2i/show_example5_init.jpg)
@@
![[Uncaptioned image]](extracted/2309.05950v5/t2i/failure_art_original.png)
@@
![[Uncaptioned image]](extracted/2309.05950v5/t2i/failure_count_original.png)
@@
![[Uncaptioned image]](extracted/2309.05950v5/t2i/failure_cube_original.png)
@@
![[Uncaptioned image]](extracted/2309.05950v5/t2i/failure_panels_original.png)
@@
![[Uncaptioned image]](extracted/2309.05950v5/t2i/japan_original.png)
![[Uncaptioned image]](extracted/2309.05950v5/t2i/japan_initial.png)
@@
![[Uncaptioned image]](extracted/2309.05950v5/t2i/panda_original.png)
![[Uncaptioned image]](extracted/2309.05950v5/t2i/panda_initial.png)